THEOREMS ON DIFFERENTIAL INEQUALITIES FOR SECOND

ORDER FUNCTIONAL DIFFERENTIAL EQUATIONS by D.Bainov, Sofia, Bulgaria, and A.Domoshnitsky, Haifa, Israel

Abstract. In the paper necessary and sufficient conditions for negativity of Green’s functions are obtained in the form of theorems on differential inequalities. As corollaries coefficient criteria for negativity of Green’s functions of two-point boundary value problems and the periodic problems are obtained. Mathematics subject classification: 34K99. Keywords: Differential inequalities, functional differential equations.